OpenStudy (anonymous):

Actually, can somebody please help me with this? (medals and fan will be given) 4x^1/2-5=27

OpenStudy (anonymous):

it was asked to be solved

OpenStudy (anonymous):

4x^1? ok. 4x^1/2=32 4x^1=16 4x=16 x=4

OpenStudy (raden):

4x^1/2 - 5 = 27 4x^1/2 = 27 + 5 4x^1/2 = 32 x^1/2 = 32/4 x^1/2 = 8 squares both sides, get x = 8^2 = 64

OpenStudy (anonymous):

oh no, the problem was